Help!
Adult tickets for a show cost $5.50 while children's tickets cost $2.50.If there were twice as many adults as children and the total receipts for the show were $1026 how many adults were at the show?

Answer by nerdybill(5412)

Adult tickets for a show cost $5.50 while children's tickets cost $2.50.If there were twice as many adults as children and the total receipts for the show were $1026 how many adults were at the show?
.
Let C = number of children
then from:"there were twice as many adults as children"
2C = number of adults
.
5.50(2C) + 2.50C = 1026
5.50(2C) + 2.50C = 1026
11C + 2.50C = 1026
13.50C = 1026
C = 1026/13.50
C = 76
.
To calculate adult tickets:
adult tickets = 2C = 2(76) = 152 (adults at show)